### v1

# A perturbative non-Markovian treatment to low-temperature spin decoherence

Abstract
Molecular spins are promising candidates for quantum information science, leveraging coherent electronic spin states for quantum sensing and computation. However, the practical application of these systems is hindered by electronic spin decoherence, driven by interactions with nuclear spins in the molecule and the surrounding environment at low temperatures. Predicting dephasing dynamics remains a formidable challenge due to the complexity of the spin bath. In this work, we develop a non-Markovian time-convolutionless master equation to treat an electronic spin coupled to a nuclear-spin bath. By relating ab initio electronic structure parameters directly to the decoherence dynamics, we provide a framework that accounts for pure dephasing in the low-temperature limit. We apply this method to a series of molecular qubit candidates and demonstrate good agreement with experimental relaxation trends. This approach offers a computationally efficient path for the prediction of low-temperature decoherence trends in molecular spin systems.
Modern technologies in quantum information science (QIS) such as quantum computing and sensing aim to leverage properties of quantum spin systems for information storage and processing beyond classical capabilities. There is significant research interest surrounding magnetically-active molecules because of their potential for room-temperature operation and the breadth of synthetic techniques available for molecular qubit design 17, 28, 33, 36, 9. Implementation of molecular spin systems typically uses an unpaired electron spin state as the qubit in QIS or perturbations of the spin states for fine magnetic field sensing. 4, 16, 43 In any application, control and manipulation of coherent spin states are necessary for robust information processing, yet present a major challenge to the viability of many quantum technologies. In low temperature molecular spin systems, fluctuations in the magnetic field interacting with the electron spin present a significant source of pure dephasing noise. Spin-active nuclei on a molecule or in a surrounding solvent system contribute a time-dependent magnetic field that can dephase the electron spin quickly. 34, 25, 32, 2 Dynamical decoupling in the form of the Hahn-echo experiment is often employed in order to remove contributions from static inhomogeneities, isolate specific decoherence mechanisms, and elongate spin dephasing time. 19, 11, 13 The fluctuation of nearby nuclear spins contribute to the decay of the Hahn echo via dipolar coupling, 10, 12, 39 but connecting the rates of spin dephasing to microscopic parameters of the molecular spin system remains a challenge due to the complexity of the electronic structure.
Current computational methods to treat electron spin dephasing include cluster-correlation expansion (CCE) approaches, 40, 41, 44, 42, 7, 30, 31 tensor-network methods, 24 and the analytical pair-product approximation (APPA). 21, 35 Despite the breadth of different methods, there is a lack of open quantum systems (OQS) master equation (ME) treatments of the pure dephasing problem. Longitudinal relaxation, $T_{1}$, of the electron spin has been treated using microscopic Markovian MEs, 1, 26, 27, 3 and the analogous relationship for $T_{2}$ dephasing has also been explored through phenomenological approaches. 22, 31 Notably microscopic MEs that explicitly relate molecular electronic structure properties to pure dephasing are desirable. Here we derive a non-Markovian master equation for the electron spin degrees of freedom to relate ab initio electronic structure to low-temperature decoherence trends where dephasing dominates. Specifically, we derive a time-convolutionless master equation to 2nd order perturbation theory (TCL2) with respect to the interaction Hamiltonian. We include a single pulse in our differential equation to better compare to Hahn-echo experiments. Through the TCL2 equation, we relate the hyperfine couplings computed with electronic structure and nuclear dipolar couplings to the coherence in the electron reduced density matrix. We obtain a result that treats the electron spin dephasing mechanism in an OQS framework with electronic structure and find that we correctly predict dephasing trends in a series of molecular qubit candidates.
We consider an electron spin doublet interacting with two spin-$\frac{1}{2}$ nuclei. We start with the spin Hamiltonian, invoke the secular approximation, and assume a large magnetic field limit such that we are in the pure dephasing regime of spin decoherence. The electron spin Hamiltonian is,
$\hat{H}_{e}=\omega_{e}\hat{S}^{z}$ (1)
where $\omega_{e}$ is the Larmor frequency and $\hat{S}^{z}$ is the Pauli-$z$ operator. The nuclear spin Hamiltonian for spins 1 and 2 is given by,
$\hat{H}_{n}=(\omega_{1}+\frac{1}{2}A_{1})\hat{I}_{1}^{z}+(\omega_{2}+\frac{1}{2}A_{2})\hat{I}_{2}^{z}+b_{12}\big(\hat{I}_{1}^{z}\hat{I}_{2}^{z}+\frac{1}{4}(\hat{I}_{1}^{+}\hat{I}_{2}^{-}+\hat{I}_{1}^{-}\hat{I}_{2}^{+})\big)$
where $\omega_{1,2}$ is the Larmor frequency, $\hat{I}_{1,2}$ are the nuclear spin operators, $A_{1}$ and $A_{2}$ are the $zz$ components of the hyperfine tensors, and $b_{12}$ is the nuclear spin-spin dipolar coupling constant.
The interaction between the electron spin and two nuclear spins is given by,
$\hat{H}_{en}(t)=\hat{S}^{z}(t)\otimes\sum_{k}A_{k}\hat{I}_{k}^{z}(t),$ (2)
where $\hat{I}_{k}^{z}(t)$ evolves in time in the interaction picture according to,
$\hat{I}_{k}^{z}(t)=e^{i\hat{H}_{n}t}\hat{I}_{k}^{z}e^{-i\hat{H}_{n}t},$ (3)
and $\hat{S}^{z}(t)$,
$\hat{S}^{z}(t)=h(t,t_{p})\cdot e^{i\hat{H}_{e}t}\hat{S}^{z}e^{-i\hat{H}_{e}t}$ (4)
where we have included the Heaviside step function modified to range in value from $1$ to 1-1, $h(t,t_{p})=(-2\Theta(t-t_{p})+1)$, to incorporate the effects of an applied pulse at time $t_{p}$.
The TCL is formally derived as a series expansion of the interaction Hamiltonian, but is often taken to only a few orders. In the limit that the magnetic field is strong, the secular spin Hamiltonian conserves the total $z$-component of the spin-angular momentum. We choose an unpolarized initial state such that all odd-numbered terms in the series expansion vanish. The lowest-order term required is the second-order term,
$\frac{d}{dt}\mathcal{P}\hat{\rho}(t)=\mathcal{K}_{2}(t)\mathcal{P}\hat{\rho}(t),$ (5)
where $\hat{\rho}(t)$ is the total density matrix of the composite system and environment, $\mathcal{P}$ is a projection operator that projects the density matrix onto the states of interest, and $\mathcal{K}_{2}(t)$ is the 2nd order TCL generator,
$\mathcal{K}_{2}(t)=\int_{0}^{t}dt_{1}\mathcal{P}\mathcal{L}(t)\mathcal{L}(t_{1})\mathcal{P},$ (6)
where $\mathcal{L}(t)$ is the Liouvillian, $\mathcal{L}(t)=-i[\hat{H}_{I}(t),\hat{\rho}(t)]$. The integral in Eq. 6 deconvolutes the density matrix in time to yield a time-local ME for $\hat{\rho}(t)$ 6. The projection operator is often chosen to project onto separable system and environment states; however, here the entanglement between subspaces is important for the dynamics. Instead, we use a correlated projection operator,
$\mathcal{P}\hat{\rho}=\sum_{m}\textrm{Tr}_{B}[\Pi_{m}\hat{\rho}]\otimes\frac{1}{N_{m}}\Pi_{m},$ (7)
where $\Pi_{m}$ is the $m^{\textrm{th}}$ eigenstate of $\hat{I}^{z}_{tot}$ and $N_{m}$ is the degeneracy. 14, 15 Because we restrict ourselves to the unpolarized state, the sum reduces to the $m=0$ state. These considerations simplify the TCL equation,
$\frac{d\hat{\rho}(t)}{dt}=$ $\bigg(2\hat{S}^{z}\hat{\rho}\hat{S}^{z}-\frac{\hat{\rho}}{2}\bigg)\int_{0}^{t}dt_{1}h(t_{1},\frac{t}{2})\langle A_{1}\hat{I}_{1}^{z}(t)A_{2}\hat{I}_{2}^{z}(t_{1})\rangle.$ (8)
The off-diagonal element of the electron reduced density matrix, or the coherence, can be found analytically,
$\rho_{e}^{01}(t)=\rho_{e}^{01}(0)e^{-W(t)},$ (9)
where
$W(t)=\int_{0}^{t}dt\int_{0}^{t}dt_{1}h(t_{1},\frac{t}{2})\langle A_{1}\hat{I}_{1}^{z}(t)A_{2}\hat{I}_{2}^{z}(t_{1})\rangle.$ (10)
We integrate to yield,
$W_{12}(t)=\bigg(\frac{2b_{12}\Delta_{12}}{b_{12}^{2}+\Delta_{12}^{2}}\bigg)^{2}\big(1-\textrm{cos}(\frac{t}{4}\sqrt{\Delta_{12}^{2}+b_{12}^{2}})\big),$ (11)
where $\Delta_{12}=A_{1}-A_{2}$. This solution is a 2nd order perturbative approach to pure dephasing of an electron interacting with two nuclear spins indexed by $1$ and $2$ in a Hahn-echo experiment. Further details of this process are shown in the Supporting Information (SI).
To extend this approach to larger systems while retaining the direct connection to molecular parameters, we split the nuclear spin correlation function into pairs. 12 This yields a factorized solution for many nuclear spins,
$\rho_{e}^{01}(t)=\rho_{e}^{01}(0)e^{-\sum_{kl}W_{kl}(t)}.$ (12)
Each nuclear spin pair $k,l$ contributes to the echo decay with an amplitude,
$\alpha_{kl}^{2}=\bigg(\frac{2b_{kl}\Delta_{kl}}{b_{kl}^{2}+\Delta_{kl}^{2}}\bigg)^{2},$ (13)
and frequency,
$f_{kl}=\frac{1}{4}\sqrt{\Delta_{kl}^{2}+b_{kl}^{2}}.$ (14)
We find that the amplitude, $\alpha_{kl}^{2}$, reaches a maximum of 1.0 when $|b_{kl}|=|\Delta_{kl}|$, and vanishes when either $b_{kl}$ or $|\Delta_{kl}|=0$ or in the limit that $|\Delta_{kl}|\rightarrow\infty$, consistent with previous results. 21, 35
We compare the dynamics of our TCL2 equation to numerically exact simulations of one electron interacting with two nuclear spins by computing the fidelity between the two density matrices,
$\mathcal{F}(\hat{\rho},\hat{\sigma})=\bigg(\textrm{Tr}[\sqrt{\sqrt{\hat{\rho}}\hat{\sigma}\sqrt{\hat{\rho}}}]\bigg)^{2}.$ (15)
The coherence of the electron spin periodically returns to a maximum value of 1 after a time interval $\tau=\frac{1}{f_{12}}$, and reaches a minimum value at $\tau/2$. Figure 1 shows the fidelity computed at $t=\tau/2$ and $t=\tau$ to evaluate the error in modulation depth and frequency, respectively. Details on the numerically exact simulations and explicit echo profiles are shown in the SI.
Figure 1: Fidelity between the TCL2 and numerically exact density matrices as a function of $\alpha_{12}^{2}$. The left side of the figure represents when $|b_{12}|<|\Delta_{12}|$ and the right side when $|b_{12}|>|\Delta_{12}|$.
The fidelity at $\tau$ remains at 1 for all values of $\alpha_{12}^{2}$, indicating that our TCL2 equation correctly predicts the frequency, $f_{12}$, of the coherence. The fidelity at $\tau/2$ decreases as $\alpha_{12}^{2}$ increases, approaching a minimum of 0.46 when $\alpha_{12}^{2}$ reaches a maximum of 1.0. Importantly, the computed coherence at $\tau/2$ from the TCL2 equation is always larger than the numerically exact density matrix; thus, the fidelity at $\tau/2$ shows that our TCL2 equation consistently underestimates the modulation depth of the coherence. Despite the deviations from the numerically exact value of the coherence, $\alpha_{12}^{2}$ does not overestimate the modulation depth, making it a good predictor of the contribution a nuclear-spin pair makes to electron dephasing in the Hahn-echo experiment.
We now demonstrate the factorized TCL2 equation in Eq. 12 by analyzing a series of previously studied vanadium-oxo molecules, [VO(C3H6S2)2]2-, [VO(C5H6S4)2]2-, [VO(C7H6S6)2]2- and [VO(C9H6S8)2]2-, labeled as V1, V2, V3 and V4, respectively. 18 The full molecular structures are shown in the SI. From V1 to V4 the spin active hydrogens become progressively further from the electron spin located on the vanadium metal center. We compute the hyperfine couplings between the unpaired electron on the vanadium ion and the distal hydrogen spins using density functional theory (DFT), and compute the dynamics of the electron spin coherence using Eq. 12.
Figure 2: Echo decay for four different vanadium-oxo molecules using the factorized TCL2 approach.
The dynamics of the electronic state in these complexes are shown in Figure 2. Here we do not see complete decoherence but rather so-called residual coherence, which is expected for small nuclear spin baths as demonstrated in previous computational studies. 7 Experimental data taken on these molecules shows that coherence decays to zero within the time scale shown in Fig. 218 To better represent the experiment we embed the molecule in a bath of hydrogen spins, simulating hydrogens present on nearby vanadium-oxo or solvent molecules. We add those contributions to Eq. 12, and show these dynamics in Figure 3.
Figure 3: Echo decay of the vanadium-oxo molecules interacting with a bath of hydrogen spins.
In the presence of additional hydrogen spins, we see that the echo decays completely, with V1 taking the longest to decay to zero, consistent with experimental data. 18
Notably in the above data, we did not consider decoherence contributions from heteronuclear spin pairs. Molecular qubit candidates often include transition metal ions or deuterated solvent molecules with $I>\frac{1}{2}$ nuclei, which can interact with hydrogen spins via dipolar interactions. We can use the TCL2 framework to evaluate the impact of heteronuclear spin pair fluctuations on the echo decay.
II Example Nuclei $W(t)$
11 2D 10-14
$3/2$ 63Cu 10-11
$5/2$ 55Mn 10-12
$7/2$ 51V 10-12
Table 1: Contributions to the echo decay for spin pairs with one spin $I>\frac{1}{2}$.
We solve Eq. 10 for several possible nuclear spins with different $I$ interacting with a spin-$\frac{1}{2}$ nucleus. For different spin states, we compute the heteronuclear correlation function and solve Eq. 10 to obtain the heteronuclear pair contribution to spin decoherence. Table 1 shows the order of magnitude of $W(t)$, which can be inserted into Eq. 12 to simulate the damping of the dynamics from a heteronuclear spin pair. For the heteronuclear spin pairs considered, $W(t)$ is vanishingly small, indicating that spin-spin interactions between heteronuclear spin pairs, such as vanadium and hydrogen, do not contribute to the echo decay for spin-spin electron dephasing within the secular spin Hamiltonian at high magnetic field.
We have presented a non-Markovian ME approach to pure dephasing in molecular spin systems at low temperature in a large magnetic field. We analyze the performance of the TCL2 equation against numerically exact simulations for a three-spin system and we find good agreement. Specifically, the frequency of oscillations generated by our TCL2 approach agrees with the numerically exact solution for all parameter regimes tested. The TCL2 approach consistently underestimates the modulation depth of the electron coherence, but provides a good prediction of the contribution of each nuclear spin pair to the decoherence. We then extend the TCL2 technique to molecular spin systems using a factorization approach of the spin-spin correlation functions, computing hyperfine coupling tensors with electronic structure calculations. Here we see trends in the molecular series which agree with previous computational studies, where the smaller molecule retains more coherence than larger molecules. To agree with previous experimental studies, we embed the molecules in a bath of hydrogen spins and simulate complete decoherence of the electronic spin state within the experimentally observed timescale. In these molecular studies, we neglect any electron decoherence contributions from heteronuclear spin-spin interactions. To justify this approximation, we compute $W(t)$ for several nuclear spins with $I>\frac{1}{2}$ commonly present in molecular qubits, and show that these contributions are vanishingly small.
The promise of molecular QIS relies on controlling and leveraging the diversity of molecular conformations and electronic structures. Relating the ab initio electronic properties to the time-dependent behavior of molecules in magnetic fields at low temperatures is essential for robust understanding of electron-nuclear spin decoherence in molecules. Our results directly relate molecular properties, including hyperfine tensors and nuclear dipolar coupling, to low-temperature decoherence dynamics in molecular spin systems. Importantly, the approach requires only a single electronic structure calculation per species. This is important in cases where the molecule cannot be described by DFT, but requires more costly techniques to simulate the correlated electronic structure. Furthermore, our technique will allow for greater understanding of how electron delocalization and correlation amplify or diminish nuclear-spin decoherence effects.
1 Computational Details
We obtain hyperfine tensors for V1 - V4 from electronic structure calculations using the ORCA 6 quantum chemistry package. 29 Following previous work, we optimize the geometries using density functional theory with B3LYP and density fitting, with a def2-triple-$\zeta$ polarized basis for hydrogen, sulfur, carbon, and oxygen and a core-polarized CP(PPP) basis for vanadium. 7, 8, 20, 37, 38, 23, 5 The vanadium-oxygen bond is oriented to lie along the $z$-axis, and we compute hyperfine tensors from the optimized geometries and use the $zz$ component in Eq. 12 for each hydrogen spin. We include contributions to the hyperfine tensors from the dipolar coupling between the electron and nuclear spins along with the isotropic Fermi contact term. We include a bath of random hydrogen spins for comparison to experimental results. We constrain the average distance from the electron to the added hydrogen spins to be greater than 11 Å, and compute the additional interactions from the point-dipole approximation. The randomized coordinates of the extra hydrogen spins were generated once and used for each of V1 - V4.
2 Associated Content
2.1 Data Availability Statement
The data and codes underlying this study are openly available at https://doi.org/10.5281/zenodo.18246098.
2.2 Supporting Information
Details on the derivation of the TCL equation, the spin Hamiltonian, the error analysis, heteronuclear spin-pairs, and the molecular systems of interest.
3 Acknowledgements
KHM acknowledges start-up funding from the University of Minnesota.